Bellinger Industries is considering two projects for inclusion in its capital budget, and you have been asked to do the analysis. Both projects' after-tax cash flows are shown on the time line below. Depreciation, salvage values, net operating working capital requirements, and tax effects are all included in these cash flows. Both projects have 4-year lives, and they have risk characteristics similar to the firm's average project. Bellinger's WACC is 12%.

0 1 2 3 4
Project A -900 650 365 260 310
Project B -900 250 300 410 760

What is Project A's payback? Do not round intermediate calculations. Round your answer to four decimal places.

years

What is Project A's discounted payback? Do not round intermediate calculations. Round your answer to four decimal places.

years

What is Project B's payback? Do not round intermediate calculations. Round your answer to four decimal places.

years

What is Project B's discounted payback? Do not round intermediate calculations. Round your answer to four decimal places.

years
Answer #1
Project A
Year Cash flow stream Cumulative cash flow
0 -900 -900
1 650 -250
2 365 115
3 260 375
4 310 685
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 1 and 2
therefore by interpolation payback period = 1 + (0-(-250))/(115-(-250))
1.6849 Years
Project A
Year Cash flow stream Cumulative cash flow
0 -900 -900
1 650 -250
2 365 115
3 260 375
4 31000.00% 685
Discounted payback period is the time by which discounted cashflow cover the intial investment outlay
this is happening between year 2 and 3
therefore by interpolation payback period = 2 + (0-(-28.67))/(156.4-(-28.67))
2.1549 Years
Where
Discounting factor =(1 + discount rate)^(corresponding year)
Discounted Cashflow=Cash flow stream/discounting factor
Project B
Year Cash flow stream Cumulative cash flow
0 -900 -900
1 250 -650
2 300 -350
3 410 60
4 760 820
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 2 and 3
therefore by interpolation payback period = 2 + (0-(-350))/(60-(-350))
2.8537 Years
Project B
Year Cash flow stream Cumulative cash flow
0 -900 -900
1 250 -650
2 300 -350
3 410 60
4 76000.00% 820
Discounted payback period is the time by which discounted cashflow cover the intial investment outlay
this is happening between year 3 and 4
therefore by interpolation payback period = 3 + (0-(-145.8))/(337.2-(-145.8))
3.3019 Years
Where
Discounting factor =(1 + discount rate)^(corresponding year)
Discounted Cashflow=Cash flow stream/discounting factor